# Anderson localization for the multi-frequency quasi-periodic CMV matrices and quantum walks

### v1

@@ Abstract
In this paper, we establish Anderson localization for the CMV matrices with multi-frequency analytic quasi-periodic Verblunsky coefficients in the regime of the positive Lyapunov exponent. As an application, we further derive the Anderson localization for the multi-frequency analytic quasi-periodic quantum walks. We extend the results of Wang and Damanik [25] for one-frequency quasi-periodic CMV matrices to multi-frequency case.
@@
1 Introduction and the statement of main result
@@
The problem of Anderson localization (AL) for CMV matrices and quantum walks has received attention recently. Zhu [29] showed both Anderson localization and strong dynamical localization for random CMV matrices. Wang-Damanik [25] formulated AL for quasiperiodic CMV matrices and quantum walks whose Verblunsky coefficients are defined by shift. Cedzich-Werner [6] obtained AL for a class of electric quantum walks and skew-shift CMV matrices. Lin-Piao-Guo [21] proved AL for CMV matrices with Verblunsky coefficients defined by skew-shift. They extended the localization results for Schrödinger operators in [2] and [4] to CMV matrices respectively. In the sequel, Lin-Guo-Piao [20] proved AL for CMV matrices whose Verblunsky coefficients are given by a hyperbolic toral automorphism. They extended the localization results for Schrödinger operators in [2] and [4] to CMV matrices.
To the best of our knowledge, the localization problem of the multi-frequency quasi-periodic CMV matrices still remains unknown. This is precisely the issue that the present paper focuses on. We are strongly motivated by the papers [7, 2, 3, 17, 28, 15, 16] whose are about AL and related spectral analysis for the multi-frequency quasi-periodic Schrödinger operators. Generally speaking, in the spectral theory of quasi-periodic Schrödinger operators, we observe that the analysis of AL for multi-frequency operators is more complicated than the one-frequency case. The most obvious obstruction is the elimination of double resonances along the orbit; see the references [1, 7]. Our study indicates that the CMV matrix situation is also similar. Indeed, there is a close relation between the spectral theories of CMV operators and one-dimensional Schrödinger operators. The detailed review for the relationship one can refer to the references [21, 25, 22, 23]. But it is necessary to emphasize that working out the CMV analog of a Schrödinger result is not straightforward.
For the one-frequency quasi-periodic CMV matrices, Wang-Damanik [25] got the elimination of double resonances along the orbit directly by using the frequency estimate [2, Lemma 6.1]. However, to our knowledge, there is no such estimate for multi-frequency CMV matrices. For this issue, we apply the semialgebraic sets arguments [1, Corollary 9.7 and Lemma 9.9]. Therefore, one of the key steps is to construct a proper semialgebraic set and do the corresponding decomposition. The other key steps include mainly the avalanche principle (AP) and the large deviation theorem (LDT).
@@
For recent work on other spectral properties of CMV matrices, such as the regularity of Lyapunov exponent, absolutely continuous spectrum, etc., there are also fruitful results and one can consult the interesting papers [10, 11, 12, 19, 26, 27].
The main result of this paper can be stated as follows:
Theorem 1.1.
The multi-frequency quasi-periodic CMV matrices $\mathcal{E}$ whose Verblunsky coefficients are defined by a shift display Anderson localization in the regime of positive Lyapunov exponents for a fixed phase vector $x_{0}$ and almost all frequency vectors satisfying the standard Diophantine condition; that is, $\mathcal{E}(x_{0})$ has pure point spectrum with exponentially decaying eigenfunctions, where $\mathcal{E}(x_{0})$ denotes the extended CMV matrix with phase $x_{0}$.
@@
The rest of the present paper is organized as follows: we provide a quick introduction to the orthogonal polynomials on the unit circle (OPUC) and the settings in Section 2. Section 3 contains some basic tools which are helpful to the following study. Section 4 addresses the proof of the main theorem. Finally, we show AL for the analytic multi-frequency quasi-periodic quantum walks in Section 5.
2 Preliminaries
@@
Let us recall some preliminaries of OPUC that can be seen for more details in [22]. Let $\mathbb{D}=\{z:|z|<1\}$ be the open unit disk in $\mathbb{C}$ and $\mu$ be a nontrivial probability measure on $\partial\mathbb{D}=\{z:|z|=1\}$. Therefore, the functions $1,z,z^{2},\ldots$ are linearly independent in the Hilbert space $L^{2}(\partial\mathbb{D},d\mu)$. Let $\Phi_{n}(z)$ be the monic orthogonal polynomials; that is, $\Phi_{n}(z)=P_{n}[z^{n}]$, $P_{n}\equiv$ projection onto $\{1,z,z^{2},\ldots,z^{n-1}\}^{\bot}$ in $L^{2}(\partial\mathbb{D},d\mu)$. Then the orthonormal polynomials are
@@
For any polynomial $Q_{n}(z)$ of degree $n$, one can define its Szegő dual $Q_{n}^{*}(z)$ by
@@
Then the Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$ ($\alpha_{n}\in\mathbb{D}$) obey the equation
@@
which is called Szegő recursion.
Orthonormalizing $1,z,z^{-1},z^{2},z^{-2},\cdots$, one can get a CMV basis $\{\chi_{j}\}_{j=0}^{\infty}$ of $L^{2}(\partial\mathbb{D},d\mu)$, and the matrix representation of multiplication by $z$ relative to the CMV basis gives rise to the CMV matrix $\mathcal{C}$,
@@
$\mathcal{C}=\left[\begin{array}[]{ccccccc}\overline{\alpha}_{0}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ \rho_{0}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&\ddots&\ddots&\ddots\end{array}\right],$
@@
$\mathcal{E}=\left[\begin{array}[]{cccccccc}\ddots&\ddots&\ddots&&&&&\\ &-\overline{\alpha}_{0}\alpha_{-1}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ &-\rho_{0}\alpha_{-1}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &&\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&&\ddots&\ddots&\ddots\end{array}\right],$
which is a special five-diagonal doubly-infinite unitary matrix in the standard basis of $L^{2}(\partial\mathbb{D},d\mu)$ according to [22, Subsection 4.5] and [23, Subsection 10.5].
In this paper, we consider a sequence of Verblunsky coefficients which are generated by an analytic function $\alpha(\cdot):\mathbb{T}^{d}\rightarrow\mathbb{D}$, $\alpha_{n}(x)=\alpha(T^{n}x)=\alpha(x+n\omega)$, where $T$ is an invertible map of the form $Tx=x+\omega$, $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$, $x,\omega\in\mathbb{T}^{d}$ are phase and frequency, respectively. Moreover, the frequency vector $\omega\in\mathbb{T}^{d}$ satisfies the standard Diophantine condition
@@
[…]
@@
$\left[\begin{array}[]{c}\varphi_{n+1}\\ \varphi_{n+1}^{*}\end{array}\right]=S(\omega,z;x+n\omega)\left[\begin{array}[]{c}\varphi_{n}\\ \varphi_{n}^{*}\end{array}\right],$
@@
$S(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}z&-\overline{\alpha}(x)\\ -\alpha(x)z&1\end{array}\right].$
@@
$M(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\alpha}(x)}{\sqrt{z}}\\ -\alpha(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right]\in\mathbb{S}\mathbb{U}(1,1).$
@@
$M_{n}(\omega,z;x)=\prod_{j=n-1}^{0}M(\omega,z;x+j\omega)$ (2.6)
@@
exists. This is called the Lyapunov exponent. Throughout this paper, we let $\gamma$ be the lower bound of the Lyapunov exponent. On the other hand, Furstenberg-Kesten theorem indicates that the limit also exists for $a.e.\;x$:
@@ 3 Basic tools
In this section, we require some useful lemmas which could have applications in the following research. To begin with, we introduce some useful notations.
Definition 1.
[5, Definition 2.3] For any positive numbers $a,b$ the notation $a\lesssim b$ means $Ca\leq b$ for some constant $C>0$. By $a\ll b$ we mean that the constant $C$ is very large. If both $a\lesssim b$ and $a\gtrsim b$, then we write $a\asymp b$.
@@ 3.1 Avalanche principle
[14, Proposition 2.2] Let $A_{1},\ldots,A_{m}$ be a sequence of arbitrary unimodular $2\times 2$-matrices. Suppose that
@@ 3.2 Large deviation theorem
[14, Proposition 9.1] Let $d$ be a positive integer. Suppose $u:D(0,2)^{d}\rightarrow[-1,1]$ is subharmonic in each variable; i.e., $z_{1}\rightarrow u(z_{1},z_{2},\ldots,z_{d})$ is subharmonic for any choice of $(z_{2},\ldots,z_{d})\in D(0,2)^{d-1}$ and similarly for each of the other variables. Assume furthermore that for some $n\geq 1$
@@
[…]
@@
Suppose $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. There exist $\sigma=\sigma(p,q)$, $\tau=\tau(p,q)$, $\sigma,\tau\in(0,1)$, $C_{0}=C_{0}(p,q,h)$ such that for $n\geq 1$ one has that
@@
Then $u_{n}$ is a continuous subharmonic function which is bounded by $1$ in $D(0,1)^{d}$. In addition, $u_{n}$ satisfies the conditions in the above lemma. Thus, (3.6) is an immediate consequence of (3.5). ∎

### v3

@@ Abstract
In this paper we prove Anderson localization for multi-frequency quasi-periodic extended CMV matrices with analytic Verblunsky coefficients in the regime of positive Lyapunov exponents. By constructing a suitable semialgebraic set and combining the Avalanche Principle with a Large Deviation Theorem, we overcome the key obstruction of eliminating double resonances along the orbit, where multi-frequency potentials introduce significant challenges compared to the single-frequency case. As a direct application, we establish Anderson localization for corresponding analytic multi-frequency quasi-periodic quantum walks via unitary equivalence.
@@
1 Introduction and Preliminaries
1.1 Research Background
@@
The study of Anderson localization (AL) for CMV matrices and quantum walks has garnered significant recent interest. Zhu [33] [^1] In the paper [33], Zhu also corrects some of the existing CMV related issues in the literature. Our present paper fully considers these corrections. established both AL and strong dynamical localization for the random CMV matrices. In the quasi-periodic setting, Wang and Damanik [29] formulated AL for CMV matrices and quantum walks with Verblunsky coefficients defined by a simple shift. Cedzich and Werner [7] obtained AL for a class of electric quantum walks and skew-shift CMV matrices. Subsequently, Lin et al. [24] proved AL for CMV matrices with Verblunsky coefficients defined by a skew-shift, extending localization results for Schrödinger operators from [4]. Following this, Lin et al. [25] also demonstrated AL for CMV matrices where the Verblunsky coefficients are generated by a hyperbolic toral automorphism.
To the best of our knowledge, the localization problem for multi-frequency quasi-periodic CMV matrices remains open. Addressing this gap is the primary focus of the present paper. Our motivation stems strongly from the extensive body of work on AL and related spectral analysis for multi-frequency quasi-periodic Schrödinger operators [11, 2, 3, 21, 32, 19, 20]. A fundamental observation in the spectral theory of quasi-periodic operators is that the analysis of AL for multi-frequency potentials presents considerably greater complexity than the one-frequency case. A principal obstruction lies in eliminating double resonances along the orbit, a well-documented challenge [1, 11]. Our investigation reveals that the situation for CMV matrices is analogous. While close connections exist between the spectral theories of CMV operators and Schrödinger operators, translating results from the Schrödinger setting to the CMV case is often non-trivial.
This non-triviality is particularly evident in the multi-frequency context. For the one-frequency quasi-periodic CMV matrices, Wang and Damanik [29] were able to directly eliminate double resonances using a frequency estimate [2, Lemma 6.1]. However, no analogous estimate is currently available for the multi-frequency CMV setting. To overcome this hurdle, we employ powerful tools from the theory of semialgebraic sets, specifically [1, Corollary 9.7 and Lemma 9.9]. Consequently, a key step in our approach is the construction of a suitable semialgebraic set and its decomposition. This strategy, combined with the application of the Avalanche Principle (AP) and a Large Deviation Theorem (LDT), forms the core of our proof.
@@
For completeness, we note that significant progress has also been made on other spectral properties of CMV matrices, such as the regularity of the Lyapunov exponent and the nature of the absolutely continuous spectrum; see, for instance, [14, 15, 16, 23, 30, 31].
@@
The rest of the present paper is organized as follows: Section 1 provides a quick introduction to the orthogonal polynomials on the unit circle (OPUC) and the settings of CMV matrices, Szegö cocycle, and Lyapunov exponents. Section 2 contains some basic tools (including the Avalanche Principle, Large Deviation Theorem, estimates for the Lyapunov exponent, Green’s function, Poisson’s formula, and semialgebraic sets) which are helpful to the following study. Section 3 addresses the proof of the main theorem (Theorem 1.1). Finally, Section 4 applies these results to derive AL for the analytic multi-frequency quasi-periodic quantum walks, utilizing the Gesztesy-Zinchenko cocycle framework.
1.2 CMV Matrices and Key Concepts
Let us recall some preliminaries of OPUC that can be seen for more details in Simon’s monograph [26].
@@ 1.2 CMV Matrices and Key Concepts
Let $\mathbb{D}=\{z:|z|<1\}$ be the open unit disk in $\mathbb{C}$ and $\mu$ be a nontrivial probability measure on $\partial\mathbb{D}=\{z:|z|=1\}$. Then the functions $1,z,z^{2},\ldots$ are linearly independent in the Hilbert space $L^{2}(\partial\mathbb{D},d\mu)$. Let $\Phi_{n}(z)$ be the monic orthogonal polynomials, that is, $\Phi_{n}(z)=P_{n}[z^{n}]$, $P_{n}\equiv$ projection onto $\{1,z,z^{2},\ldots,z^{n-1}\}^{\bot}$ in $L^{2}(\partial\mathbb{D},d\mu)$. Then the orthonormal polynomials are
@@
For any polynomial $Q_{n}(z)$ of degree $n$, its Szegő dual $Q_{n}^{*}(z)$ is defined by
@@
Then Szegő recursion is
@@
where the parameters $\{\alpha_{n}\}_{n=0}^{\infty}$ in $\mathbb{D}$ are called the Verblunsky coefficients.
Orthonormalizing $1,z,z^{-1},z^{2},z^{-2},\cdots$, yields a CMV basis $\{\chi_{j}\}_{j=0}^{\infty}$ of $L^{2}(\partial\mathbb{D},d\mu)$, and the matrix representation of multiplication by $z$ relative to the CMV basis gives rise to the CMV matrix $\mathcal{C}$,
@@
$\mathcal{C}=\left(\begin{array}[]{ccccccc}\overline{\alpha}_{0}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ \rho_{0}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&\ddots&\ddots&\ddots\end{array}\right),$
@@
$\mathcal{E}=\left(\begin{array}[]{cccccccc}\ddots&\ddots&\ddots&&&&&\\ &-\overline{\alpha}_{0}\alpha_{-1}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ &-\rho_{0}\alpha_{-1}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &&\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&&\ddots&\ddots&\ddots\end{array}\right),$
which is a special five-diagonal doubly-infinite unitary matrix in the standard basis of $L^{2}(\partial\mathbb{D},d\mu)$ according to [26, Subsection 4.5] and [27, Subsection 10.5].
1.3 Szegő Cocycle and Lyapunov Exponent
In this paper, we consider the extended CMV matrix with Verblunsky coefficients generated by an analytic function $\alpha(\cdot):\mathbb{T}^{d}\rightarrow\mathbb{D}$, $d>1$, $\alpha_{n}(x)=\alpha(T^{n}x)=\alpha(x+n\omega)$, where $T:\mathbb{T}^{d}\rightarrow\mathbb{T}^{d}$ is the shift map of the form $x=x+\omega, 𝕋:=/\mathbb{T}$$\mathbb{R}$$\mathbb{Z}$, $x,\omega\in\mathbb{T}^{d}$ are phase and frequency, respectively. Moreover, the frequency vector $\omega\in\mathbb{T}^{d}$ satisfies the standard Diophantine condition
@@ 1.3 Szegő Cocycle and Lyapunov Exponent
for all nonzero $k\in\mathbb{Z}^{d}$, where $p>0$, $q>d$ are some constants, $\|\cdot\|$ denotes the distance to the nearest integer and $|\cdot|$ stands for the sup-norm on $\mathbb{Z}^{d}$ ($|k|=|k_{1}|+|k_{2}|+\cdots+|k_{d}|$, $k_{i}$ is the $i$-th element of the vector $k$).
For the sake of convenience, the set of $\omega$ satisfying the standard Diophantine condition (1.2) is denoted by $\mathbb{T}^{d}(p,q)\subset\mathbb{T}^{d}$.
@@
$\left(\begin{array}[]{c}\varphi_{n+1}\\ \varphi_{n+1}^{*}\end{array}\right)=S(\omega,z;x+n\omega)\left(\begin{array}[]{c}\varphi_{n}\\ \varphi_{n}^{*}\end{array}\right),$
@@
$S(\omega,z;x)=\frac{1}{\rho(x)}\left(\begin{array}[]{cc}z&-\overline{\alpha}(x)\\ -\alpha(x)z&1\end{array}\right).$
@@
$M(\omega,z;x)=\frac{1}{\rho(x)}\left(\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\alpha}(x)}{\sqrt{z}}\\ -\alpha(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right)\in\mathbb{S}\mathbb{U}(1,1).$
@@
$M_{n}(\omega,z;x)=\prod_{j=n-1}^{0}M(\omega,z;x+j\omega).$ (1.6)
@@
exists. This is called the Lyapunov exponent. Throughout this paper, we let $\gamma$ be the lower bound of the Lyapunov exponent. On the other hand, Furstenberg-Kesten theorem indicates that the limit also exists for a.e.xa.e.\;x:
@@
Definition 1.1 (Generalized eigenvalues and generalized eigenfunctions).
We call $z$ a generalized eigenvalue of the extended CMV matrix $\mathcal{E}$ , if there exists a nonzero, polynomially bounded function $u$ such that $\mathcal{E}=zu$ where $u=\{u(n)\}_{n\in\mathbb{Z}}$. We call $u$ a generalized eigenfunction.
Definition 1.2 (Anderson localization).
We say that extended CMV matrix $\mathcal{E}$ exhibits Anderson localization on $\mathcal{I}\subset\partial\mathbb{D}$ , if $\mathcal{E}$ has only pure point spectrum in $\mathcal{I}$ and for every eigenvalue its eigenfunction $u$ decay exponentially in $n$.
1.4 Main Results
Our main results are following theorem and applications to quantum walks.
Theorem 1.1.
Let $\mathcal{I_{\alpha}}\subset\partial\mathbb{D}$ be a compact spectral interval. Consider the extended CMV matrices $\mathcal{E}(x)$ with Verblunsky coefficients $\alpha_{n}(x)=\alpha(x+n\omega)$, where $\alpha:\mathbb{T}^{d}\to\mathbb{D}$ is analytic. Assume that for every $z\in\mathcal{I_{\alpha}}$ and every $\omega\in\mathbb{T}^{d}(p,q)$, there exists a constant $\gamma>0$ such that the Lyapunov exponent satisfies $L(\omega,z)>\gamma>0$. Then for fixed $x_{0}$ and a.e. $\omega\in\mathbb{T}^{d}(p,q)$, $\mathcal{E}(x_{0})$ exhibits Anderson localization on $\mathcal{I_{\alpha}}$.
This theorem has applications to a class of quantum walks; see Theorem 4.2.
@@ 2 Basic Tools
In this section, we require some useful lemmas which could have applications in the following research. To begin with, we introduce some useful notations[5, Definition 2.3].
For any positive numbers $a,b$ the notation $a\lesssim b$ means $Ca\leq b$ for some constant $C>0$. By $a\ll b$ we mean that the constant $C$ is very large. If both $a\lesssim b$ and $a\gtrsim b$, then we write $a\asymp b$.
@@ 2.1 Avalanche Principle
[18, Proposition 2.2] Let $A_{1},\ldots,A_{m}$ be a sequence of arbitrary unimodular $2\times 2$-matrices. Suppose that
@@ 2.2 Large Deviation Theorem
[18, Proposition 9.1] Let $d$ be a positive integer. Suppose $u:D(0,2)^{d}\rightarrow[-1,1]$ is subharmonic in each variable; i.e., $z_{1}\rightarrow u(z_{1},z_{2},\ldots,z_{d})$ is subharmonic for any choice of $(z_{2},\ldots,z_{d})\in D(0,2)^{d-1}$ and similarly for each of the other variables. Assume furthermore that for some $n\geq 1$
@@
Here $\langle u\rangle=\int_{\mathbb{T}^{d}}u(\theta)d\theta$. If $d=2$ then the range is $0<\tau<\frac{1}{3}-\varepsilon_{2}$ and $\sigma=\frac{1}{3}-\tau-\varepsilon_{2}$ where $\varepsilon_{2}\rightarrow 0$ as $\varepsilon_{1}\rightarrow 0$.
@@
Let $\mathcal{I_{\alpha}}\subset\partial\mathbb{D}$ be a compact spectral interval. Suppose that for all $\omega\in\mathbb{T}^{d}(p,q)$ and $z\in\mathcal{I_{\alpha}}\subset\mathbb{D}$, $L(\omega,z)>\gamma>0$. There exist $\sigma=\sigma(p,q)$, $\tau=\tau(p,q)$, $\sigma,\tau\in(0,1)$, $C_{0}=C_{0}(p,q,h)$ such that for $n\geq 1$ one has that
@@
Then $u_{n}$ is a continuous subharmonic function which is bounded by 11 in $D(0,1)^{d}$. In addition, $u_{n}$ satisfies the conditions in the above lemma. Thus, (2.6) is an immediate consequence of (2.5). ∎